# Supplementary Information

### v1

@@ Results
To eliminate extreme outliers, we implemented the modified IQR outlier detection rule (described in Eq. (8)) to identify data points that lay unusually far from their respective cluster centers. This procedure eliminated roughly $2\%$ of the data points, resulting in a retention of $\sim 98\%$, and ensured that subsequent calculations of any cluster-based statistics were not skewed by outliers, while keeping the cluster assignments unchanged. For the grouped data, clustering revealed three states with lifetimes of $0.74\pm 0.15\,\text{ns}$, $2.01\pm 0.47\,\text{ns}$, and $1.8\pm 0.1\,\text{ns}$, respectively, with an average weighted lifetime of $1.73\pm 0.55\,\text{ns}$. There are $\frac{N(N-1)}{2}$ possible reversible pairwise switching pathways for $N$ clusters. The dynamics of the three pathways between the three clusters for the Alexa data are summarized in Table 1, indicating that Alexa tends to switch preferably from state 1 to state 2 ($R_{12}>>1$), from state 3 to state 1 ($R_{13}<1$), and from state 3 to state 2 ($R_{23}<<1$).
Clusters (i,j) $k_{\text{{i}}\text{{j}}}\,(\text{s}^{-1})$ $k_{\text{{ji}}}\,(\text{s}^{-1})$ $f\,(\text{s}^{-1})$ $R_{\textit{ij}}$
(1, 2) 6.06 1.25 2.08 4.84
(1, 3) 4.29 7.99 5.59 0.54
(2, 3) 0.134 1.06 0.23 0.13
Table 1: Switching rates ($k_{ij}$), switching frequencies ($f$), and switching ratios ($R_{ij}$) between the relevant clustered states of Alexa, using Eqs. (5)-(7).
@@
To test the robustness of the clustering analysis to statistically small datasets, we considered a dataset comprising only 30 QDs, measured for 30 s each. QD 605 consists of a CdSe core surrounded by a ZnS shell. When the shell is sufficiently thin, these QDs, which belong to type II-IV QDs, typically exhibit bimodal blinking behavior (palstra2021python, efros2016origin). A representative intensity trace is shown in Figure 4A, illustrating the relatively fast photoblinking dynamics of this sample, although, unlike the Alexa data, the intensity levels could be well-resolved by the change-point algorithm. The lifetime–intensity distributions for the ungrouped and grouped QD data are shown in Figure 4B and C, respectively. The grouping effectively reduces the noise in the data, as evidenced by the markedly narrower lifetime–intensity distribution, which uncovers the underlying linear correlation between the two parameters. The significant denoising facilitated the subsequent data clustering. The corresponding BIC score plots show a shallow gradient from two clusters and higher for both data sets (Figs. 4 D and E), suggesting that any of two to eight clusters would be an appropriate choice, although two is the preferred number as it occurs after the first steepest gradient. The results of a two-cluster model are shown in Figures 4F and G for the ungrouped and grouped data, respectively, with ellipses indicating data assignment per cluster shown at a confidence of $0.85$.
@@
We implemented the modified IQR outlier detection rule (Eq. (8)) on the grouped data prior to calculating the switching rates and frequencies. Approximately $4\%$ of the data points were removed, resulting in a retention of $\sim 96\%$ of the original data. This step ensured that extreme outliers were disregarded so that they would not skew the computed dynamics. The lifetimes corresponding to the two cluster centers of the grouped data are $2.2\pm 1.2\,\text{ns}$ and $9.5\pm 3.3\,\text{ns}$, with a weighted average of $7.0\pm 4.6\,\text{ns}$. The switching rates between the two states were $k_{12}=1.68\,\text{s}^{-1}$ and $k_{21}=3.04\,\text{s}^{-1}$, with a switching frequency of $f=2.16\,\text{s}^{-1}$, indicating that switching from state 2 to state 1 is more likely, i.e., state 1 is the more stable state, in agreement with its higher population density.
@@
Figure 4: Clustering results for the statistically small QD 605 data with $\sim 30$ particles. A. Representative example intensity trace showing 40-ms binned data (grey) and resolved intensity levels (green). B and C. Lifetime–intensity distribution for the ungrouped (B) and grouped (C), respectively. D and E. Corresponding BIC score plots. F and G. Clustering both the ungrouped and grouped datasets using two clusters, with cluster centers indicated with black crosses, and ellipses (at a confidence of $0.85)$ indicating the assignment of data points to each cluster.
LHCII
The main light-harvesting complex of plants, LHCII, was chosen to test the capability of clustering a multichromophoric system that switches between different physiologically significant intensity–lifetime states and to address the debate about the actual number and identity of distinct states (schlau2015single, kruger2011fluorescence, kruger2011fluorescence2). Specifically, some studies suggest a multistate system that displays exclusively a linear correlation between intensity and lifetime, with some intermediate intensity states between the strong and weak emitting states (kruger2011fluorescence, kruger2011fluorescence2), while some other studies suggest the existence of a third, non-linear state in addition to the strongly and weakly emitting states (schlau2015single).
An example intensity trace for the grouped LHCII data is shown in Figure 5A, which illustrates photoblinking between unquenched, quenched, and intermediate intensity states. The blinking behavior is considerably less erratic than that observed for Alexa and QD 605. The ungrouped and grouped lifetime–intensity distributions are shown in Figures 5B and C, respectively, displaying a clear linear relationship that becomes more prominent when reducing the noise by means of grouping (Fig. 5C). Notably, the data shows no indication of the existence of the third, long-lifetime and low-intensity population that was reported in Ref. (schlau2015single). The BIC score plot from the clustering analysis indicates three clusters as the statistically most likely outcome for both the ungrouped (Fig. 5D) and grouped (Fig. 5E) data, and the results of clustering both datasets, using three clusters, are shown in Figures 5F and G, respectively. The cluster centers confirm the linear correlation between the lifetime and intensity for both the ungrouped and grouped data. It is clear that grouping makes the clustering more accurate, leading to tighter and better-defined clusters.
Figure 5: Clustering outcome for the LHCII data, comparing the ungrouped (left) and grouped (right) results for a system of 102 particles. A. Example intensity trace with 40-ms binned data (grey) and the resolved levels (green). B and C. Lifetime–intensity distributions for the ungrouped (B) and grouped (C) LHCII data. D and E. BIC scores plots for the ungrouped and grouped data, respectively. F and G. Corresponding clustering results using three clusters, with cluster centers indicated by black crosses, and ellipses drawn at a confidence of $0.85$.
The clustering revealed three states with lifetimes of $0.56\pm 0.18\,\text{ns}$, $2.1\pm 0.7\,\text{ns}$, and $3.56\pm 0.12\,\text{ns}$, respectively, with an average weighted lifetime of $2.7\,\pm 1.3\text{ ns}$. IQR filtering removed approximately $2\%$ of extreme outlier data points, retaining $\sim 98\%$ of the dataset. This preserved cluster assignments while ensuring no outliers contributed to the dynamics between clusters. The switching parameters, determined using Eqs. (5)-(7), are summarized in Table 2, indicating that LHCII tends to switch more regularly from state 2 to state 1 ($R_{12}<1$), from state 1 to state 3 ($R_{13}>1$), and from state 2 to state 3 ($R_{23}>1$). This points to a dynamic equilibrium between the three states that is shifted toward the unquenched state (state 3, representing the light-harvesting function of the complex) and away from the short-living, partially quenched state 2.
Clusters (i,j) $k_{\text{{i}}\text{{j}}}\,(\text{s}^{-1})$ $k_{\text{{ji}}}\,(\text{s}^{-1})$ $f\,(\text{s}^{-1})$ $R_{\textit{ij}}$
(1, 2) 0.21 0.28 0.24 0.75
(1, 3) 0.31 0.11 0.16 2.8
(2, 3) 0.39 0.11 0.17 3.54
Table 2: Switching rates ($k_{ij}$), switching frequencies ($f$), and switching ratios ($R_{ij}$) for the LHCII three-clustered grouped data, using Eqs. (5), (7).
Phycobilisome (PB)
PB from Synechocystis PCC6803 was chosen as the last case study due to its size, complexity, and biological significance. Comprising a few dozen polypeptides and a total of 396 pigments (dominguez2022structures), it can be considered the ultimate example of a large, multichromophoric system of biological importance that absorbs visible light. This photosynthetic light-harvesting complex exhibits a heterogeneous excitation energy landscape (cupellini2024energetic, kruger2019role), which results in considerable variations in its fluorescence emission, in terms of both energy and intensity (gwizdala2016controlling). Another property that puts our clustering analysis to the test is the instability of PB under in vitro conditions. Specifically, when isolated, this complex readily loses some of its subunits, invariably resulting in variations in absorption cross-section and sometimes in its fluorescence properties (gwizdala2016controlling). To address the consequent variations in fluorescence intensity, each complex’s intensity trace was normalized such that its maximum resolved intensity corresponded to the global maximum resolved intensity among all intensity traces, ensuring that each intensity trace was comparable to the next. This allowed direct comparison of intensity levels across complexes and considerably denoised the lifetime–intensity distributions. This normalization is one of the recently implemented functionalities in the Full SMS toolbox (botha2024advanced).
After the normalization and subsequent grouping, the PB data were clustered. A representative example intensity trace for the grouped PB data is shown in Figure 6A, which displays photoblinking between the relevant (normalized) intensity states. The intensity switching frequency of this sample was relatively low due to the use of a very low excitation power. The lifetime–intensity distribution of the ungrouped (Fig. 6B) and grouped (Fig. 6C) data shows significant heterogeneity, with several well-defined dense regions, which become more pronounced after grouping. The grouping again significantly denoises the data, producing better-defined data boundaries, which facilitates the clustering. Our clustering protocol was applied to both the ungrouped and grouped PB data, where for the ungrouped data, the BIC score plot (Fig. S3A) revealed five clusters as the optimal choice, whereas for the grouped data (Fig. 6D), three clusters correspond to a clear local minimum. The clustering results for the ungrouped data, using five clusters, are shown in Figure S3B, while a three-cluster model applied to the grouped data is depicted in Figure 6E. In both cases, no linear trend is evident; however, the cluster centers align well with the centers of visibly dense data regions.
For both the ungrouped and grouped data, the second local minimum corresponds to eight clusters (see Figs. S3A and 6D). Applying an eight-cluster model to the grouped data gives rise to considerable overlap between the confidence ellipses (Fig. S4), strongly suggesting overfitting. However, eight clusters are potentially a physically relevant choice for this complex (vide infra).
The IQR outlier detection rule (Eq. (8)) gave rise to $\sim 97\%$ data retention, and the three clustered states have lifetimes of $0.96\pm 0.23\,\text{ns}$, $1.5\pm 0.24\,\text{ns}$, and $1.57\pm 0.19\,\text{ns}$, respectively, with an average weighted lifetime of $1.62\,\pm 0.22$ ns. The dynamics between the clusters are summarized in Table 3, which reveals that the system tends to preferably switch from state 1 to state 2 ($R_{12}>>1$), from state 1 to 3 ($R_{13}>>1$), and from state 3 to 2 ($R_{23}<1$), indicating that state 2 is the most stable state and state 1 is the least stable.
Figure 6: Results of the clustering protocol applied to the grouped PB data with 101 measured particles. A. Example intensity trace of a PB particle, showing 40-ms binned data in grey, and the grouped intensity levels overlaid in green. B and C. Ungrouped and grouped PB lifetime–intensity distribution. D. BIC score plot for the grouped PB data. E. Corresponding clustered results for the grouped data, using three clusters, with cluster centers indicated by black crosses. Ellipses illustrate which data points are most likely associated with which cluster, at a confidence of $0.85$.
Clusters (i,j) $k_{\text{{i}}\text{{j}}}\,(\text{s}^{-1})$ $k_{\text{{ji}}}\,(\text{s}^{-1})$ $f\,(\text{s}^{-1})$ $R_{\textit{ij}}$
(1, 2) 0.075 0.008 0.014 9.38
(1, 3) 0.014 0.003 0.005 4.67
(2, 3) 0.004 0.009 0.006 0.44
Table 3: Switching rates ($k_{ij}$), switching frequencies ($f$), and switching ratios ($R_{ij}$) between the relevant clustered states of PB, using Eqs. (5)-(7)).
@@ Discussion
Owing to the high sensitivity of SMS, any sources of heterogeneity, caused by inherent experimental noise, system composition, non-optimized experimental conditions, or data fitting uncertainties, can significantly distort the data distributions by introducing broadening. This is especially true for fluorescence intensity data, as it is generally strongly dependent on the excitation probability of the sample (kruger2011fluorescence), unlike the fluorescence lifetime. On the other hand, the resolved lifetime usually has a greater uncertainty than the resolved intensity, leading to additional broadening. Data broadening most often arises from noise and is much less likely from heterogeneity in physically meaningful states.
@@
Skewed or unevenly distributed populations can bias clustering, causing cluster centers to be shifted away from where they might be expected through visual inspection. To address this, the grouping algorithm, implemented in the Full SMS software, helps to denoise the data and reveal any dominant underlying relationships between experimental parameters, thereby tightening distributions for interpretation. The effect is evident when comparing, for example, Figures 3C and 3D, where, in the latter case, a cleaner and more structured distribution is observed. Grouping also emphasizes the linear trend between lifetime and intensity in many SMS experiments (schlau2015single, botha2024advanced, gwizdala2016controlling, weston2002measuring, tinnefeld2001photophysical). By denoising and tightening the data, grouping makes subsequent clustering easier, more robust, and more reliable.
It is important to note that the clustering algorithm does not assess whether the clusters correspond to physically meaningful states or not; it simply partitions the data based on localized high-density regions. The clustering protocol fits the data with a mixture of Gaussian distributions, while the BIC score determines the optimal number of clusters by balancing fitness with complexity. In practice, this means that the algorithm favors simpler models unless the addition of more parameters would provide a substantial improvement to the fit. Notice that when performing clustering, the confidence ellipses representing relevant cluster regions may overlap. This is expected since GMM is a probabilistic method whose Gaussian components have indefinitely extending tails, unlike hard-clustering algorithms with strict boundaries (such as K-means). As a result, multiple clusters can contribute non-zero probabilities to the same data point, which is reflected by the overlapping ellipses. Nevertheless, each data point is assigned to the cluster for which it has the highest posterior probability, that is, the cluster to which it most likely belongs.
The BIC score curve does not necessarily display a monotonically decreasing trend, and therefore often exhibits several local minima. Although the global minimum suggests the most accurate statistical fit, it may lead to overfitting by assigning redundant cluster centers to a single broad state. Conversely, there is no objective justification for selecting any other local minimum, and the general assumption that a system can be modeled by a simple two-state model may be a crude oversimplification. To address this, we imposed the condition of considering the first relevant local minimum, following Occam’s razor. This approach describes the data objectively and sufficiently. Figure 3 illustrates this well. The BIC plot for the grouped data (Fig. 3F) displays multiple local minima. The first minimum, corresponding to three clusters, yields cluster centers that visually agree with the dense regions in the lifetime–intensity distributions (Fig. 3H). The second minimum, which corresponds to five clusters (Fig. S2), most likely produces overfitting since multiple cluster centers are now assigned to a single broad cluster (cluster 2 in Fig. 3H) with considerable overlap between the confidence ellipses. Nevertheless, it is still up to the user’s discretion to determine which of these states are physically significant. Caution should be exercised when certain assumptions are made about the number of relevant clusters, as this may lead to user bias, unless this decision is based on well-established knowledge about the system. The transitional dynamics between states can help to identify which clusters are physically relevant and which are not. For example, if a system with five identified clusters includes two that rarely share dynamics with other populations, they are likely artifacts, and the system may have been overfitted. In contrast, clusters that exhibit strong dynamic interactivity with one another are more likely to represent physically meaningful states.
@@
To evaluate this reasoning in a controlled manner, we benchmarked the clustering protocol using simulated datasets where the number and nature of each state are known a priori, thereby allowing direct comparison of the algorithm’s output to ground truth. The protocol was successful in clustering well-defined two- and four-state datasets (Figs. 1 and S1) with cluster centers aligning well with the centers of the densest regions of the distributions.
For the perturbed two-state dataset, the lifetime–intensity distribution (Fig. 2B) is smeared due to polarization and dipole orientation effects. Neither of the two candidate solutions in the BIC score plot assigned a data cluster near 2500 cps because its density is significantly lower than at 500 cps (cf. Fig. 2F). As a result, the algorithm—which penalizes unnecessary complexity—accepts two clusters as a sufficient choice to describe the two high-density data populations, while also suggesting three clusters as a more optimal solution, with the third cluster lying within the smeared-out region but closer to the highest-density population. This behavior of the clustering illustrates how the BIC inherently balances fit accuracy and model simplicity, accepting a simpler solution than what visual inspection may suggest, highlighting the potential bias and inaccuracy of visual approaches.
@@
For Alexa, the ungrouped data (Fig. 3C) appear more spread out with no obvious trend, whereas the grouped data (Fig. 3D) reveals the underlying trends due to substantial denoising. The large intensity spread in the grouped data corresponding to a lifetime of $\sim 1.8-2.0$ ns can be explained by Fig. 2, i.e., a strong broadening in the intensity distribution of the bright state is caused by the heterogeneous distribution of molecular orientations relative to the circularly polarized excitation light, causing large variations in the relative absorption cross-sections of the molecules. The absence of broadening down to the background level can be explained by user bias in the data selection: molecules that are oriented away from the focal plane interact only weakly with the incident light and, therefore, appear to be significantly dimmer than those oriented near-parallel to the focal plane. These dim signals are easily masked by the background noise and, as a result, those molecules are unlikely to be selected for further study. This skews the overall intensity distribution to brighter states, as is the case for Alexa (Fig. 3D).
The authors in Ref. (wang2013lifetime) report the lifetime of Alexa 647 to be $1.13\,\text{ns}$, which is longer than the bulk lifetime of 1 ns due to the use of glycerol. In our study, the sample was spin-coated and measured in the presence of oxygen. The average weighted single-molecule lifetime in our dataset ($1.73\pm 0.55\,\text{ns}$) deviates from that of previous studies, although this prolonged lifetime is not likely caused by the effects of different environmental conditions, but instead a result of the data grouping by the Full SMS software. Figures 3C and D indicate that the average lifetime of the quenched state (state 1) increased from $\sim 0.35$ ns for the ungrouped data (Fig. 3C) to $\sim 0.75$ ns for the grouped data (Fig. 3D). This indicates that most of the short-lifetime data in the ungrouped dataset was combined with longer-lifetime data during the grouping procedure, likely to limit the number of short events that lack statistical weight or are prone to fitting errors. This unavoidable artifact of data grouping can be limited by choosing another BIC value during the grouping analysis to ensure that a short-lifetime population is chosen as a distinct state.
Since Alexa switched more frequently from state 3 into states 1 and 2 (Table 1), states 2 and 3 may represent a single physically broadened state; however, previous studies (lin2015quantifying, yang2024electrochemically, herdly2023benchmarking)) suggested that Alexa may exhibit multiple distinct fluorescing states. These observations demonstrate the capability of our clustering protocol to successfully identify and separate physically meaningful subpopulations, confirming its robustness for complex datasets.

### v2

@@ Results
To eliminate extreme outliers, we implemented the modified IQR outlier detection rule (described in Eq. (12)) to identify data points that lay unusually far from their respective cluster centers. This procedure eliminated roughly $2\%$ of the data points, resulting in a retention of $\sim 97\%$, and ensured that subsequent calculations of any cluster-based statistics were not skewed by outliers, while keeping the cluster assignments unchanged. For the grouped data, clustering revealed two states with lifetimes of $0.75\pm 0.17\,\text{ns}$ and $1.94\pm 0.35\,\text{ns}$, respectively, with an average weighted lifetime of $1.73\pm 0.55\,\text{ns}$. There are $\frac{N(N-1)}{2}$ possible reversible pairwise switching pathways for $N$ clusters. The switching rates between the two states were $k_{12}=9.87\,\text{s}^{-1}$ and $k_{21}=1.99\,\text{s}^{-1}$, with a switching frequency of $f=3.31\,\text{s}^{-1}$, and switching ratio of $R_{12}>>1$ indicating that switching from state 1 to state 2 is more likely, i.e., state 2 is the more stable state, in agreement with its higher population density.
@@
To test the robustness of the clustering analysis to statistically small datasets, we considered a dataset comprising only 30 QDs, measured for 30 s each. QD 605 consists of a CdSe core surrounded by a ZnS shell. When the shell is sufficiently thin, these QDs, which belong to type II-IV QDs, typically exhibit bimodal blinking behavior (38, 42). A representative intensity trace is shown in Figure 6A, illustrating the relatively fast photoblinking dynamics of this sample, although, unlike the Alexa data, the intensity levels could be well-resolved by the change-point algorithm. The lifetime–intensity distributions for the ungrouped and grouped QD data are shown in Figure 6B and C, respectively. The grouping effectively reduces the noise in the data, as evidenced by the markedly narrower lifetime–intensity distribution, which uncovers the underlying linear correlation between the two parameters. The significant denoising facilitated the subsequent data clustering. The IC score plots for the ungrouped and grouped data are shown in Figures 6D and E, respectively. For this dataset, it is again evident that a global minimum solution, which lies even beyond ten clusters, will yield over-partitioning; hence, we instead employ our candidate cluster identification method. For the ungrouped data, the BIC and AIC curves show a shallow gradient beyond two clusters, while ICL deviates, rising after two clusters. Universally, the first steepest drop (with the greatest IC gain) occurs at two clusters, yielding $\langle D^{2}\rangle=1.89$ and FO = 0.13, indicating that two clusters are the most statistically significant choice. While AIC and BIC suggest eight clusters as a secondary option, the auxiliary metrics of $\langle D^{2}\rangle=1.6$ and FO = 0.08 deviate too much from their optimal values, indicating that the two-cluster solution provides the best optimal choice between compactness and coverage of the data, for the candidate numbers of clusters. For the grouped dataset, AIC and BIC again show a near-identical trend with the strongest elbow at two clusters, while ICL suggests additional possible partitions at five, seven, and ten clusters. However, the two-cluster solution (with $\langle D^{2}\rangle=1.89$ and FO = 0.11) remains the most meaningful partition of the grouped data, representing physically relevant states while avoiding overfitting. Figures 6F and G show the resulting two-cluster solutions for the ungrouped and grouped data, respectively.
@@
We implemented the modified IQR outlier detection rule (Eq. (12)) on the grouped data prior to calculating the switching rates and frequencies. Approximately $4\%$ of the data points were removed, resulting in a retention of $\sim 96\%$ of the original data. This step ensured that extreme outliers were disregarded so that they would not skew the computed dynamics. The lifetimes corresponding to the two cluster centers of the grouped data are $2.2\pm 1.2\,\text{ns}$ and $9.5\pm 3.3\,\text{ns}$, with a weighted average of $7.0\pm 4.6\,\text{ns}$. The switching rates between the two states were $k_{12}=1.68\,\text{s}^{-1}$ and $k_{21}=3.04\,\text{s}^{-1}$, with a switching frequency of $f=2.16\,\text{s}^{-1}$, and with a switching ratio of $R_{12}<1$ indicating that switching from state 2 to state 1 is more likely, i.e., state 1 is the more stable state, in agreement with its higher population density.
@@
Figure 6: Clustering results for the statistically small QD 605 data with $\sim 30$ particles. A. Representative example intensity trace showing 40-ms binned data (grey) and resolved intensity levels (green), with the dashed blue line indicating the background intensity level. B and C. Lifetime–intensity distribution for the ungrouped (B) and grouped (C), respectively. D and E. Corresponding IC score plots. F and G. Clustering both the ungrouped and grouped datasets using two clusters, with cluster centers indicated by black crosses, and ellipses (at a confidence of $0.85)$ indicating the assignment of data points to each cluster.
@@ Discussion
Owing to the high sensitivity of SMS, even modest experimental noise, system heterogeneity, suboptimal experimental conditions, or fitting uncertainties can substantially broaden measured data distributions. This is especially true for fluorescence intensity data, as it is generally strongly dependent on the excitation probability of the sample (62). On the other hand, the resolved lifetime usually has a greater uncertainty than the resolved intensity, leading to additional broadening along the lifetime axis. Consequently, such broadening is most often dominated by noise and measurement or analysis uncertainty rather than genuine heterogeneity in the underlying physical states.
@@
Skewed or unevenly distributed populations can bias clustering, causing cluster centers to be shifted away from where they might be expected through visual inspection. To address this, the grouping algorithm, implemented in the Full SMS software, helps to denoise the data and reveal any dominant underlying relationships between experimental parameters, thereby tightening distributions for interpretation. The effect is evident when comparing, for example, Figures 5C and 5D, where, in the latter case, a cleaner and more structured distribution is observed. Grouping also emphasizes the linear trend between lifetime and intensity in many SMS experiments (18, 55, 63, 64, 65). By denoising and tightening the data, grouping makes subsequent clustering easier, more robust, and more reliable.
Comparison of alternative clustering methods on well-defined simulated datasets (Figs. S2 and S4) highlights the advantages of the GMM-based protocol. SMS data are naturally elliptical with non-zero covariance between intensity and lifetime, a structure that GMMs inherently capture through full covariance estimation, whereas DPGMM and K-means either depend sensitively on priors or assume spherical clusters. DPGMM can recover the underlying populations but is highly sensitive to prior specification, which can lead to low-occupancy “ghost” clusters as observed in Figures S2A and S4B. K-means, being a hard-assignment method, relies on subjective elbow-point selection and does not provide probabilistic assignments, making it prone to misclassification unless the number of clusters is manually set (see Fig. S4). Both methods are also computationally slower than GMM with IC-based selection. Overall, GMM with IC metrics provides more reliable, objective, and efficient clustering of SMS datasets without requiring prior knowledge of the number of populations.
It is important to note that the clustering algorithm does not assess whether the clusters correspond to physically meaningful states or not; it simply partitions the data based on localized high-density regions. The protocol models the data as a Gaussian mixture, with the optimal cluster number determined using multiple IC metrics (AIC, BIC, and ICL), which balance goodness-of-fit, model complexity, and classification uncertainty. For SMS datasets in general, primarily due to their intrinsic noise and measurement uncertainty, IC score plots often exhibit several candidate elbow points rather than a single monotonic trend. While the global minimum provides the statistically best fit, in experimental datasets, it may correspond to unrealistically high cluster numbers that result in over-partitioning and lack physical interpretability (see, for example, Figs. 5E–F and 6D–E). The clustering algorithm removes a substantial portion of subjectivity by automatically determining optimal candidates based on relevant IC gain and steep gradient descent in the plots. To further mitigate ambiguity, auxiliary cluster-quality metrics—average cluster tightness and FO—are evaluated alongside the IC scores. Cluster tightness quantifies compactness relative to dimensionality, while FO assesses whether the confidence regions adequately capture the data spread. This combined framework ensures that selected solutions are statistically consistent and interpretable, even when IC trends alone do not uniquely identify an optimum. In the Alexa dataset (Figs. 5E and F), multiple candidate solutions are present; however, only the selected solutions simultaneously maintain reasonable tightness and data coverage, preventing over-segmentation and yielding physically interpretable results. While AIC and BIC generally exhibit similar trends, ICL often diverges beyond the most prominent elbow points. This suggests that, for practical purposes, AIC and/or BIC alone may be sufficient in future studies.
@@
The simulated benchmarks were constructed as a sequence of controlled stress tests, progressing from idealized, well-separated systems to increasingly overlap-dominated and experimentally distorted regimes, in order to identify both the operational range and the intrinsic limits of the clustering protocol. Because the ground truth is known in all simulations, these tests allow direct distinction between algorithmic performance and fundamental limits imposed by data geometry. For clean two- and four-state systems with equal occupancies and minimal overlap (Figs. 1 and S1), the protocol consistently recovers the correct number of clusters, with ideal compactness ($\langle D^{2}\rangle=2$), FO values near the expected 0.05, and unity ARI. These results establish a baseline: when populations are statistically well-separable, the IC metrics, confidence ellipses, and auxiliary quality measures are mutually consistent, and recovery is unambiguous. Importantly, increasing the number of well-defined states does not degrade performance, confirming that the protocol scales correctly when overlap remains low. Polarization and dipole-orientation effects introduce the first nontrivial distortion by broadening intensity distributions without adding new physical states (Fig. 2). Importantly, an additional cluster may arise that does not correspond to a true physical state but arises from the asymmetric intensity distribution created by polarization effects. This demonstrates that immediate, polarization-based broadening contributes to apparent cluster splitting, and as shown in subsequent multi-state simulations, such effects can compound with diffuse populations and overlap to produce unexpected or “nonphysical” clustering behavior.
The influence of geometric overlap was isolated by adding additional states at fixed occupancy while systematically varying only their distribution widths (Fig. 3). These results indicate that as the state overlap increases, clustering progressively becomes more ambiguous due to reduced separability between underlying states. Importantly, the clustering method remains non-restrictive: higher-cluster solutions (e.g., five clusters) remain accessible if desired, but are automatically flagged as statistically weak, thereby avoiding subjective selection while preserving user flexibility. The loss of ground-truth recoverability at SD = $0.3\times$ mean is also physically expected, as such extreme broadening is unlikely under typical experimental conditions. However, significant overlap may occur in experimental datasets comprising a larger number of states with more moderate broadening, although experimentally induced broadening can often be mitigated through grouping strategies, as demonstrated in Fig. 5D, which effectively tightens populations.
Varying state occupancies at fixed distribution widths primarily alters relative population density rather than geometric separability, as observed in Fig. 4. For small to moderate occupancy variations, the preferred cluster number and the intrinsic quality of candidate solutions—quantified by cluster tightness and coverage—remain essentially unchanged, indicating that clustering performance is largely insensitive to modest probability shifts when overlap is fixed. However, for sufficiently large occupancy imbalances, a qualitative change emerges: low-probability states are effectively down-weighted and are no longer resolved individually, but instead become coupled into a single cluster, while the high-occupancy, well-defined states remain accurately recovered. Unlike distribution broadening, which directly increases overlap and limits recoverability across all states, changes in occupancy influence clustering only once extreme probability disparities are reached. Overall, the absolute probability of occurrence has a far weaker impact on clustering performance than the distribution width and the resulting geometric overlap.
Across all simulation classes, including those with alternative blinking kinetics, compactness and FO remain stable while ARI degrades only when overlap becomes dominant, confirming that clustering performance is controlled by the geometric structure of the observable space rather than by dwell-time statistics. Furthermore, benchmarking across multiple blinking models confirms that variations in on- and off-state dwell-time distributions have minimal impact on the inferred clustering, as the algorithm is governed primarily by the geometric structure of the lifetime–-intensity distributions rather than by the detailed temporal kinetics of state switching. Together, these results demonstrate that the proposed clustering framework is both robust and broadly applicable under diverse experimental conditions.
@@
For Alexa, the ungrouped data (Fig. 5C) appear more spread out with no obvious trend, whereas the grouped data (Fig. 5D) reveals the underlying trends due to substantial denoising. The large intensity spread in the grouped data corresponding to a lifetime of $\sim 1.8-2.0$ ns can be explained by Figure 2, i.e., a strong broadening in the intensity distribution of the bright state is caused by the heterogeneous distribution of molecular orientations relative to the circularly polarized excitation light, causing large variations in the relative absorption cross-sections of the molecules. The absence of broadening down to the background level can be explained by user bias in the data selection: molecules that are oriented away from the focal plane interact only weakly with the incident light and, therefore, appear to be significantly dimmer than those oriented near-parallel to the focal plane. These dim signals are easily masked by the background noise and, as a result, those molecules are unlikely to be selected for further study. This skews the overall intensity distribution to brighter states, as is the case for Alexa (Fig. 5D).
The authors in Ref. (23) report the lifetime of Alexa 647 to be $1.13\,\text{ns}$, which is longer than the bulk lifetime of 1 ns due to the use of glycerol. In our study, the sample was spin-coated and measured in the presence of oxygen. The average weighted SM lifetime in our dataset ($1.73\pm 0.55\,\text{ns}$) deviates from that of previous studies, although this prolonged lifetime is not likely caused by the effects of different environmental conditions, but instead a result of the data grouping by the Full SMS software. Figures 5C and D indicate that the average lifetime of the quenched state (state 1) increased from $\sim 0.35$ ns for the ungrouped data (Fig. 5C) to $\sim 0.75$ ns for the grouped data (Fig. 5D). This indicates that most of the short-lifetime data in the ungrouped dataset was combined with longer-lifetime data during the grouping procedure, likely to limit the number of short events that lack statistical weight or are prone to fitting errors. This unavoidable artifact of data grouping can be limited by choosing another IC value during the grouping analysis (in the Full SMS software) to ensure that a short-lifetime population is chosen as a distinct state.
State 2 represents a single physically broadened state; however, previous studies (66, 67, 68) suggested that Alexa may exhibit multiple distinct fluorescing states. These additional states may be masked by the broadened population, emphasizing the importance of ensuring optimal experimental conditions. Nevertheless, the presence of multiple candidate solutions in the IC scores demonstrates that our clustering protocol can reliably identify and separate physically meaningful subpopulations, thereby allowing the user to select an alternative elbow point if prior knowledge about the expected number of states is available.